OpenStudy (anonymous):

solve the inequality of solve the inequality of /x-800/≤80

OpenStudy (binary3i):

|x-800|<=80 -80<=x-800<=80 x is in [720,880]

OpenStudy (anonymous):

whenever you have an absolute value inequality like the one you have, you need to solve these two inequalities: x-800 <= 80 AND x-800 >=-80

OpenStudy (anonymous):

notice the two inequalities are joined by the conjunction AND. this will determine your solutions set to your inequality

OpenStudy (anonymous):

can you solve the first inequality: x-800 <= 80

OpenStudy (anonymous):

i got it thanks

OpenStudy (anonymous):

yw... the answer posted by binary3i is what u should get when solving these inequalities separately . ... good work... :)